By selling a fan for RS 475 a person loses 5%. To get a gain of 5%.he should sale the fan for.
Found 2 solutions by FrankM, stanbon:
Answer by FrankM(1040)

475 is 95% of the break even. We want a 5% profit or 105%
475*105/95 = 525.

Answer by stanbon(75887)

By selling a fan for RS 475 a person loses 5%. To get a gain of 5%.he should sale the fan for.
Find Initial fan value:: 0.95v = 475
Initial value = 500 RS
His lose on the 475 sale = 25 RS
-----------------
Equation
To recoup his lose and get a gain of 5% on the initial value solve for x
x = 25 + 1.05*500 = 550 RS
-----------
Cheers,
Stan H.